OpenStudy (anonymous):

What is the value of 42 – 2(3 • 5 + 1)? (1 point)

OpenStudy (anonymous):

use BODMAS rule...

OpenStudy (anonymous):

i tried that about 3 times but it didn't work help me

OpenStudy (unklerhaukus):

\[42 – \left\{2\left[(3 \times 5) + 1\right]\right\}\]

OpenStudy (anonymous):

42-2(9) 42-18 14

OpenStudy (anonymous):

42 - 2*(15 + 1) 42 - 2*(16) 42-32 10

OpenStudy (anonymous):

sorry muy mistake